OpenStudy (anonymous):

find the following integral. Use the fundamental theorem of Calculus when appropriate

OpenStudy (anonymous):

\[\int\limits_{1}^{-2}3x ^{2}dx\]

OpenStudy (anonymous):

'anti derivative' is \[x^3\] plug in -2, plug in 1, subtract the second from the first

OpenStudy (anonymous):

\[3(-2)^{2}d(-2)\]

OpenStudy (anonymous):

like this?

OpenStudy (anonymous):

should get -9

OpenStudy (anonymous):

please show work

OpenStudy (anonymous):

Satellite has done it above ^^ just do this: (-2)^3 - (1)^3 = -8 -1 = -9

OpenStudy (anonymous):

what callum said